OpenStudy (just_one_last_goodbye):

@Lucinda12345 Kevin and Randy have a jar containing 77 coins, which are all quarters or nickels. The total value of all the coins is $12.65 How many of each type of coin do they have?

OpenStudy (anonymous):

You need to open it under a new question like @just_one_last_goodbye did here for you :)

OpenStudy (anonymous):

You need to set up two equations to determine this. The first equation will represent the number of coins. Use Q to represent quarters and N to represent Nickels. So waht would the equation for the number of coins look like?

OpenStudy (just_one_last_goodbye):

Oh ok @nelsonjedi will help you ^-^ while i go help the user for health science

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

q+n=77

OpenStudy (anonymous):

Very good. Now we need an equation to represent the total amount of the coins. How would you write that equation?

OpenStudy (anonymous):

t stands for total?

OpenStudy (anonymous):

Think about how much a quarter and nickel represent in value. Total would be $ 12.65.

OpenStudy (anonymous):

so qn77=12.65?

OpenStudy (anonymous):

Let's forget about the first equation we did. That will come in later. Quarters equal 0.25 Nickels equal 0.05 The total of the number of nickles plus the number of quarters equals $ 12.65 So what would the equation look like?

OpenStudy (anonymous):

what I sometimes do is guess like 30 quarters and 47 nickles and if its too much then have more nickles and less quarters. this takes forever, and i know there is an easier algebraic way that i don't know

OpenStudy (anonymous):

i don't know much about equations

OpenStudy (anonymous):

oh i think i know!

OpenStudy (anonymous):

xQ+yN=77

OpenStudy (anonymous):

If we were looking at just quarters for an equation say how many quarters are in $1.00 and quarters equal 0.25. With Q representing the number of quarters. How would you write this as an equation?

OpenStudy (anonymous):

and 0.25x+0.5y=12.65

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

0.25 x 4 = 1.00? or is that not an equation

OpenStudy (anonymous):

0.25x + 0.05y = 12.65 Just off a decimal place. Very good. But instead of using x and y lets use q and n.

OpenStudy (anonymous):

ok so 0.25q + 0.05n = 12.65

OpenStudy (anonymous):

So we would have 0.25q +0.05n = 12.65 and from our first equation we had q+n=77..Now you need to use substitution to solve.

OpenStudy (anonymous):

so it would be 0.25q + 0.05n = 12.65 q + n = 77

OpenStudy (anonymous):

so q = 77-n

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

so 0.25(77-n) + 0.05n = 12.65

OpenStudy (anonymous):

Good..go on

OpenStudy (anonymous):

19.25 - 0.25n +0.05 n = 12.65 and then combinde like terms right?

OpenStudy (anonymous):

19.25 - 0.20n = 12.65 -0.20n = 12.65-19.25

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

-0.20n = -6.6

OpenStudy (anonymous):

0.20n = 6.6

OpenStudy (anonymous):

n = 6.6 divided by 0.20??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

n = 33 so 33 nickels and 44 quarters

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!! :)

OpenStudy (anonymous):

Very good job..and you are welcome